Let {an}∞n=1 and {bn}∞n=1 be sequences both of which diverge to −∞. Then {an +bn}∞n=1 diverge to −∞ and {an ·bn}∞n=1 diverge to +∞


1
Expert's answer
2021-02-02T12:31:23-0500

Here we have "\{a_n\}_{n\geq1}" and "\{b_n\}_{n\geq1}" both are divergent sequences, which diverge to "-\infin" .


  • "\{a_n+b_n\}_{n\geq1}"


Let us assume that "\{a_n+b_n\}_{n\geq1}" converges. Then we know by the property of arithmeticity of sequences, both "\{a_n\}_{n\geq1}" and "\{b_n\}_{n\geq1}" converge.

But we already have the fact that "\{a_n\}_{n\geq1}" and "\{b_n\}_{n\geq1}" are divergent sequences.


So, we get a contradiction.


Hence our assumption is false, i.e. "\{a_n+b_n\}_{n\geq1}" diverges.

Also, as "\{a_n\}_{n\geq1}" and "\{b_n\}_{n\geq1}" diverge to "-\infin" , by arithmetic property we have the sequence "\{a_n+b_n\}_{n\geq1}" diverges to "-\infin" .



  • "\{a_n\cdot b_n\}_{n\geq1}"


Let us assume that "\{a_n\cdot b_n\}_{n\geq1}" converges. Then we know that "\{a_n\cdot b_n\}_{n\geq1}" is bounded.

So, "\exist" c"\in\R" such that "|a_n\cdot b_n|\leq c" "\forall" n.

Since "b_n" diverges to "-\infin" , we can get "b_n<0" for all n.

So, we now have,


"|a_n|=|\frac{a_n\cdot b_n}{b_n}|\leq\frac{c}{|b_n|}" for all n.


So ,we get that "\{a_n\}_{n\geq1}" converges to 0, which is a contradiction as we already have "\{a_n\}_{n\geq1}" diverges to "-\infin" .


Hence our assumption is false, and "\{a_n\cdot b_n\}_{n\geq1}" diverges.

Also, as "\{a_n\}_{n\geq1}" and "\{b_n\}_{n\geq1}" diverge to "-\infin" , by arithmetic property we have the sequence "\{a_n\cdot b_n\}_{n\geq1}" diverges to "\infin" .